From: NASA HQ/SpaceRef/NASA Watch
Posted: Wednesday, September 14, 2005
SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by SpaceRef.com (copyright © 2005) to enhance access to related status reports and NASA activities.
All ISS systems continue to function nominally, except those noted previously or below.
The crew continued Progress offloading and cargo transfers, in the process updating the IMS (Inventory Management System).
Later, working in the Service Module (SM) and Docking Compartment (DC-1), CDR Krikalev dismantled the Russian Matryoshka payload system and prepared it for return to Earth, making sure that toxic materials in the sensors have not contaminated the worksites, by using the AK-1M air sampler and IPD Draeger tubes for ammonia. [Matryoshka automatically took radiation measurements in the SM and DC-1 docking compartment for studies of on-orbit radiation and long-term dose accumulation, using six SPD dosimeters deployed throughout the Russian segment (RS) as well as in a spherical body-simulating Matryoshka-R phantom and a human torso model outside on the SM hull, mounted there during EVA-9 on 2/27/04 and returned on EVA-14.]
Last day of the three-day transition of the onboard PCS (Portable Computer System) to software Version R9 and the ThinkPad A31p platforms, without crew involvement. [The ground today took the Standby C&C MDM (Command & Control Multiplexer/Demultiplexer) temporarily off (to Diagnostic mode), loaded it with the appropriate R9 files and reinitialized it. This committed all three C&Cs to the PCS R9, concluding the job. Ending configuration is: C&C-3 primary, C&C-2 backup, C&C-1 standby.]
To reflect the R9 upgrade in the onboard Warning books, FE/SO Phillips made appropriate SODF (Station Operations Data File) pen-and-ink changes in Caution & Warning procedures on the MPV (Manual Procedure Viewer).
With the Elektron O2 generator still off (to be repaired tomorrow), Sergei Krikalev worked on the Russian BMP harmful impurities removal system, starting the "bake-out" cycle to vacuum on absorbent bed #2 of the regenerable dual-channel filtration system. Before sleep time today, the bake-out will be terminated. Channel 1 was regenerated last Monday. [Regeneration of each of the two cartridges takes about 12 hours and is conducted only during crew awake periods. The BMP currently still uses the same vacuum vent valve for regeneration as the Elektron (the latter for venting hydrogen).]
The crewmembers each completed their portions of the weekly TVIS maintenance, i.e., inspecting their SLD (subject loading device) harnesses and checking the condition of the SPDs (subject positioning devices) and roller bearings, as well as recording time & date values.
Sergei did the daily routine maintenance of the SM's environment control & life support system (SOZh), including its toilet system (ASU) and today also the weekly routine checkup of the IP-1 airflow sensors in the various RS hatchways and FGB-to-Node tunnel, while John prepared the regular daily IMS delta /update file for automated export/import to the three IMS databases (at MCC-H, TsUP, and Baikonur).
Both crewmembers completed their regular 2.5-hr. physical exercise program on the TVIS treadmill, RED resistive machine and VELO bike with bungee cord load trainer. [Sergei s daily protocol prescribes a strict four-day microcycle exercise with 1.5 hr on the treadmill and one hour on VELO plus load trainer (today: Day 4 of a new set).]
Afterwards, John transferred the exercise data files to the MEC (Medical Equipment Computer) for downlink, as well as the daily wristband HRM (heart rate monitor) data of the RED workouts, followed by their erasure on the HRM storage medium (done six times a week).
Working off his discretionary time permitting task list, the CDR was to perform the regular daily inspection of the BIO-5 Rasteniya-2 ("Plants-2") experiment, including filling its water canister for the Lada-7 greenhouse as required.
Until the Elektron oxygen (O2) generator is brought back on line with the new BZh-8 Liquid Unit (to be installed tomorrow), cabin atmosphere is being refreshed with O2 from Progress tankage as needed to maintain proper ppO2 (oxygen partial pressure).
Over the past weekend the crew successfully completed the zero calibration on the two U.S. CSA-CP (Compound Specific Analyzer-Combustion Products) units. Now both units' CO sensors are calibrated, and recent readings indicate that there is no further concern for CO air quality.
Today's CEO (crew earth observations) photo targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, were S Mozambique vegetation/geology (general orientational obliques [in preparation for more detailed images] are requested of the Mozambique coastline for this new site. The geology of the area is unknown in detail. Vegetation types should be indicative of underlying geology. Specifically, local knowledge suggests elephant salt licks should be visible as open patches in this remote savanna forest. The open patches may permit salt bed mapping), Gulf coast (Dynamic event. There has been a special request for glint views of the flood affected areas along the Gulf coast. The crew had an opportunity, looking between popcorn cumulus, to image flood-water surfaces near the glint point [at nadir and aft in terms of the velocity vector]. Glint views reveal the margins of water bodies with great precision, even through hazy air), and Hurricane Ophelia (Dynamic event. Ophelia will be a Category 1 storm making landfall on the N Carolina coast when ISS passed overhead.)
CEO photography can be viewed and studied at the websites:
See also the website "Space Station Challenge" at:
To view the latest photos taken by the expedition 11 crew visit:
Expedition 11 Flight Crew Plans can be found at http://spaceflight.nasa.gov/station/timelines/
Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.
ISS Orbit (as of this morning, 7:21am EDT [= epoch]):
Upcoming Events (all times EDT):
Apogee height -- Mean Altitude -- Perigee height
For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see http://www.hq.nasa.gov/osf/station/viewing/issvis.html. In addition, information on International Space Station sighting opportunities can be found at http://spaceflight.nasa.gov/realdata/sightings/ on NASA's Human Spaceflight website. The current location of the International Space Station can be found at http://science.nasa.gov/temp/StationLoc.html at NASA's Marshall Space Flight Center. Additional satellite tracking resources can be found at http://www.spaceref.com/iss/tracking.html.
// end //